# Matrix Manifold Neural Networks++

### v1

@@ 1 Introduction
In recent years, deep neural networks on Riemannian manifolds have achieved impressive performance in many applications (Ganea et al., 2018; Skopek et al., 2020; Cruceru et al., 2021; Shimizu et al., 2021). The most popular neural networks in this family operate on hyperbolic spaces. Such spaces of constant sectional curvature, like spherical spaces, have the rich algebraic structure of gyrovector spaces. The theory of gyrovector spaces (Ungar, 2002; 2005; 2014) offers an elegant and powerful framework based on which natural generalizations (Ganea et al., 2018; Shimizu et al., 2021) of essential building blocks in DNNs are constructed for hyperbolic neural networks (HNNs).
Matrix manifolds such as SPD and Grassmann manifolds offer a convenient trade-off between structural richness and computational tractability (Cruceru et al., 2021; López et al., 2021). Therefore, in many applications, neural networks on matrix manifolds are attractive alternatives to their hyperbolic counterparts. However, unlike the approaches in Ganea et al. (2018); Shimizu et al. (2021), most existing approaches for building SPD and Grassmann neural networks (Dong et al., 2017; Huang & Gool, 2017; Huang et al., 2018; Nguyen et al., 2019; Brooks et al., 2019; Nguyen, 2021; Wang et al., 2021) do not provide necessary techniques and mathematical tools to generalize a broad class of DNNs to the considered manifolds.
Recently, the authors of Kim (2020); Nguyen (2022b) have shown that SPD and Grassmann manifolds have the structure of gyrovector spaces or that of nonreductive gyrovector spaces (Nguyen, 2022b) that share remarkable analogies with gyrovector spaces. The work in Nguyen & Yang (2023) takes one step forward in that direction by generalizing several notions in gyrovector spaces, e.g., the inner product and gyrodistance (Ungar, 2014) to SPD and Grassmann manifolds. This allows one to characterize certain gyroisometries of these manifolds and to construct MLR on SPD manifolds.
Although some useful notions in gyrovector spaces have been generalized to SPD and Grassmann manifolds  (Nguyen, 2022a; b; Nguyen & Yang, 2023) that set the stage for an effective way of building neural networks on these manifolds, many questions remain open. In this paper, we aim at addressing some limitations of existing works using a gyrovector space approach. Our contributions can be summarized as follows:
@@
    We propose a method for performing backpropagation with the Grassmann logarithmic map in the projector perspective (Bendokat et al., 2020) without resorting to any approximation schemes. We then show how to construct graph convolutional networks (GCNs) on Grassmann manifolds.
@@ 2.1 SPD Manifolds
The space of $n\times n$ SPD matrices, when provided with some geometric structures like a Riemannian metric, forms SPD manifold $\operatorname{Sym}_{n}^{+}$ (Arsigny et al., 2005). Data lying on SPD manifolds are commonly encountered in various domains (Huang & Gool, 2017; Brooks et al., 2019; Nguyen, 2021; Sukthanker et al., 2021; Nguyen, 2022b; a; Nguyen & Yang, 2023). In many applications, the use of Euclidean calculus on SPD manifolds often leads to unsatisfactory results (Arsigny et al., 2005). To tackle this issue, many Riemannian structures for SPD manifolds have been introduced. In this work, we focus on two widely used Riemannian metrics, i.e., Affine-Invariant (AI) (Pennec et al., 2004) and Log-Euclidean (LE) (Arsigny et al., 2005) metrics, and a recently introduced Riemannian metrics, i.e. Log-Cholesky (LC) metrics (Lin, 2019) that offer some advantages over Affine-Invariant and Log-Euclidean metrics.
@@ 2.2 Grassmann Manifolds
Grassmann manifolds $\operatorname{Gr}_{n,p}$ are the collection of linear subspaces of fixed dimension $p$ of the Euclidean space $\mathbb{R}^{n}$ (Edelman et al., 1998). Data lying on Grassmann manifolds arise naturally in many applications (Absil et al., 2007; Bendokat et al., 2020). Points on Grassmann manifolds can be represented from different perspectives (Bendokat et al., 2020). Two typical approaches use projection matrices or those with orthonormal columns. Each of them can be effective in some problems but might be inappropriate in some other contexts (Nguyen, 2022b). Although geometrical descriptions of Grassmann manifolds have been given in numerous works (Edelman et al., 1998), some computational issues remain to be addressed. For instance, the question of how to effectively perform backpropagation with the Grassmann logarithmic map in the projector perspective remains open.
@@ 2.3.1 Neural Networks on SPD Manifolds
The work in Huang & Gool (2017) introduces SPDNet with three novel layers, i.e., Bimap, LogEig, and ReEig layers that has become one of the most successful architectures in the field. In Brooks et al. (2019), the authors further improve SPDNet by developing Riemannian versions of batch normalization layers. Following these works, some works (Nguyen et al., 2019; Nguyen, 2021; Wang et al., 2021; Kobler et al., 2022; Ju & Guan, 2023) design variants of Bimap and batch normalization layers in SPD neural networks. The work in Chakraborty et al. (2020) presents a different approach based on intrinsic operations on SPD manifolds. Their proposed layers have nice theoretical properties. A common limitation of the above works is that they do not provide necessary mathematical tools for constructing many essential building blocks of DNNs on SPD manifolds. Recently, some works (Nguyen, 2022a; b; Nguyen & Yang, 2023) take a gyrovector space approach that enables natural generalizations of some building blocks of DNNs, e.g., MLR for SPD neural networks.
@@ 2.3.2 Neural Networks on Grassmann Manifolds
In Huang et al. (2018), the authors propose GrNet that explores the same rule of matrix backpropagation (Ionescu et al., 2015) as SPDNet. Some existing works (Wang & Wu, 2020; Souza et al., 2020) are also inspired by GrNet. Like their SPD counterparts, most existing Grassmann neural networks are not built upon a mathematical framework that allows one to generalize a broad class of DNNs to Grassmann manifolds. Using a gyrovector space approach, Nguyen & Yang (2023) has shown that some concepts in Euclidean spaces can be naturally extended to Grassmann manifolds.
@@ 3.1 Notation
Let $\mathcal{M}$ be a homogeneous Riemannian manifold, $T_{\mathbf{P}}\mathcal{M}$ be the tangent space of $\mathcal{M}$ at $\mathbf{P}\in\mathcal{M}$. Denote by $\exp(\mathbf{P})$ and $\log(\mathbf{P})$ the usual matrix exponential and logarithm of $\mathbf{P}$, $\operatorname{Exp}_{\mathbf{P}}(\mathbf{W})$ the exponential map at $\mathbf{P}$ that associates to a tangent vector $\mathbf{W}\in T_{\mathbf{P}}\mathcal{M}$ a point of $\mathcal{M}$, $\operatorname{Log}_{\mathbf{P}}(\mathbf{Q})$ the logarithmic map of $\mathbf{Q}\in\mathcal{M}$ at $\mathbf{P}$, $\mathcal{T}_{\mathbf{P}\rightarrow\mathbf{Q}}(\mathbf{W})$ the parallel transport of $\mathbf{W}$ from $\mathbf{P}$ to $\mathbf{Q}$ along geodesics connecting $\mathbf{P}$ and $\mathbf{Q}$. For simplicity of exposition, we will concentrate on real matrices. Denote by $\operatorname{M}_{n,m}$ the space of $n\times m$ matrices, $\operatorname{Sym}^{+}_{n}$ the space of $n\times n$ SPD matrices, $\operatorname{Sym}_{n}$ the space of $n\times n$ symmetric matrices, $\operatorname{S}^{+}_{n,p}$ the space of $n\times n$ SPSD matrices of rank $p\leq n$, $\operatorname{Gr}_{n,p}$ the $p$-dimensional subspaces of $\mathbb{R}^{n}$ in the projector perspective. For clarity of presentation, let $\widetilde{\operatorname{Gr}}_{n,p}$ be the $p$-dimensional subspaces of $\mathbb{R}^{n}$ in the ONB (orthonormal basis) perspective (Bendokat et al., 2020). For notations related to SPD manifolds, we use the letter $g\in\{ai,le,lc\}$ as a subscript (superscript) to indicate the considered Riemannian metric, unless otherwise stated. Other notations will be introduced in appropriate paragraphs. Our notations are summarized in Appendix A.
@@ 3.2 Neural Networks on SPD Manifolds
In Nguyen (2022a; b), the author has shown that SPD manifolds with Affine-Invariant, Log-Euclidean, and Log-Cholesky metrics form gyrovector spaces referred to as AI, LE, and LC gyrovector spaces, respectively. We adopt the notations in these works and consider the case where $r=1$ (see Nguyen (2022b), Definition 3.1). Let $\oplus_{ai},\oplus_{le}$, and $\oplus_{lc}$ be the binary operations in AI, LE, and LC gyrovector spaces, respectively. Let $\ominus_{ai},\ominus_{le}$, and $\ominus_{lc}$ be the inverse operations in AI, LE, and LC gyrovector spaces, respectively. These operations are given in Appendix G.
@@ 3.2.1 FC Layers in SPD Neural Networks
Our method for generalizing FC layers to the SPD manifold setting relies on a reformulation of SPD hypergyroplanes (Nguyen & Yang, 2023). We first recap the definition of SPD hypergyroplanes.
Definition 3.1 (SPD Hypergyroplanes (Nguyen & Yang, 2023)).
@@
where .,.$\langle.,.\rangle^{g}_{\mathbf{P}}$ denotes the inner product at $\mathbf{P}$ given by the considered Riemannian metric.
@@
where $\mathbf{I}_{n}$ denotes the $n\times n$ identity matrix, and .,.$\langle.,.\rangle^{g}$ is the SPD inner product in $\operatorname{Sym}_{n}^{+,g}$ (Nguyen & Yang, 2023) (see Appendix G.7 for the definition of the SPD inner product).
@@
In DNNs, an FC layer linearly transforms the input in such a way that the $k$-th dimension of the output corresponds to the signed distance from the output to the hyperplane that contains the origin and is orthonormal to the $k$-th axis of the output space. This interpretation has proven useful in generalizing FC layers to the hyperbolic setting (Shimizu et al., 2021).
@@
Let $(\mathbf{e}_{1},\ldots,\mathbf{e}_{m}),\|\mathbf{e}_{i}\|=1,i=1\ldots,m$ be an orthonormal basis of $\mathbb{R}^{m}$. Let .,.$\langle.,.\rangle^{ai}_{\mathbf{P}}$ be the Affine-Invariant metric computed at $\mathbf{P}\in\operatorname{Sym}_{m}^{+,ai}$ as
@@
Denote by $v_{(i,j)}(\mathbf{X})=\langle\ominus_{ai}\mathbf{P}_{(i,j)}\oplus_{ai}\mathbf{X},\mathbf{W}_{(i,j)}\rangle^{ai},\mathbf{P}_{(i,j)},\mathbf{W}_{(i,j)}\in\operatorname{Sym}^{+,ai}_{n},i\leq j,i,j=1,\ldots,m$. Let $\alpha=\frac{1}{m}(\sqrt{1+m\beta}-1)$. Then the output of an FC layer is computed as $\mathbf{Y}=\exp\big{(}[y_{(i,j)}]_{i,j=1}^{m}\big{)}$, where $[y_{(i,j)}]_{i,j=1}^{m}$ is the matrix having $y_{(i,j)}$ as the element at the $i$-th row and $j$-th column, and $y_{(i,j)}$ is given by
@@
As shown in Arsigny et al. (2005), a Log-Euclidean metric on $\operatorname{Sym}_{n}^{+,le}$ can be obtained from any inner product on $\operatorname{Sym}_{n}$. In this work, we consider a metric that is invariant under all similarity transformations, i.e., the metric $\langle\mathbf{W},\mathbf{V}\rangle^{le}_{\mathbf{I}_{n}}=\operatorname{Tr}(\mathbf{W}\mathbf{V})$. We have the following result.
@@ 3.2.2 Convolutional Layers in SPD Neural Networks
where $\mathbf{x}_{(i,j)}^{l}$ is a receptive field of the $l$-th input channel, $\mathbf{w}^{(l,k)}$ is the filter associated with the $l$-th input channel and the $k$-th output channel, and $b^{k}$ is the bias for the $k$-th output channel. Let $\mathbf{X}_{(i,j)}=\operatorname{concat}(\mathbf{x}_{(i,j)}^{1},\ldots,\mathbf{x}_{(i,j)}^{N_{in}})$, $\mathbf{W}_{k}=\operatorname{concat}(\mathbf{w}^{(1,k)},\ldots,\mathbf{w}^{(N_{in},k)})$, where operation $\operatorname{concat}(.)$ concatenates all of its arguments. Then Eq. (1) can be rewritten (Shimizu et al., 2021) as
@@
In Chakraborty et al. (2020), the authors design a convolution operation for SPD neural networks. However, their method is based on the concept of weighted Fréchet Mean, while ours is built upon the concepts of SPD hypergyroplane and SPD pseudo-gyrodistance from an SPD matrix to an SPD hypergyroplane (Nguyen & Yang, 2023). Also, our convolution operation can be used for dimensionality reduction, while theirs always produces an output of the same dimension as the inputs.

### v2

@@ 1 Introduction
In recent years, deep neural networks on Riemannian manifolds have achieved impressive performance in many applications (NEURIPS2018_dbab2adc; skopek2020mixedcurvature; CruceruAAAI21; shimizu2021hyperbolic). The most popular neural networks in this family operate on hyperbolic spaces. Such spaces of constant sectional curvature, like spherical spaces, have the rich algebraic structure of gyrovector spaces. The theory of gyrovector spaces (UngarBeyonEinstein02; UngarAnalyticHyp05; UngarHyperbolicNDim) offers an elegant and powerful framework based on which natural generalizations (NEURIPS2018_dbab2adc; shimizu2021hyperbolic) of essential building blocks in DNNs are constructed for hyperbolic neural networks (HNNs).
Matrix manifolds such as SPD and Grassmann manifolds offer a convenient trade-off between structural richness and computational tractability (CruceruAAAI21; FedericoGyrocalculusSPD21). Therefore, in many applications, neural networks on matrix manifolds are attractive alternatives to their hyperbolic counterparts. However, unlike the approaches in NEURIPS2018_dbab2adc; shimizu2021hyperbolic, most existing approaches for building SPD and Grassmann neural networks (DongSPD17; HuangGool17; HuangAAAI18; NguyenFG19; NguyenHandRecgCVPR19; BrooksRieBatNorm19; NguyenHandRecgICPR20; Nguyen_2021_ICCV; WangSymNet21) do not provide necessary techniques and mathematical tools to generalize a broad class of DNNs to the considered manifolds.
Recently, the authors of KimSYMGyroOrder20; NguyenNeurIPS22 have shown that SPD and Grassmann manifolds have the structure of gyrovector spaces or that of nonreductive gyrovector spaces (NguyenNeurIPS22) that share remarkable analogies with gyrovector spaces. The work in NguyenGyroMatMans23 takes one step forward in that direction by generalizing several notions in gyrovector spaces, e.g., the inner product and gyrodistance (UngarHyperbolicNDim) to SPD and Grassmann manifolds. This allows one to characterize certain gyroisometries of these manifolds and to construct MLR on SPD manifolds.
Although some useful notions in gyrovector spaces have been generalized to SPD and Grassmann manifolds  (NguyenECCV22; NguyenNeurIPS22; NguyenGyroMatMans23) that set the stage for an effective way of building neural networks on these manifolds, many questions remain open. In this paper, we aim at addressing some limitations of existing works using a gyrovector space approach. Our contributions can be summarized as follows:
@@
    We propose a method for performing backpropagation with the Grassmann logarithmic map in the projector perspective (bendokat2020grassmann) without resorting to any approximation schemes. We then show how to construct graph convolutional networks (GCNs) on Grassmann manifolds.
@@ 2.1 SPD Manifolds
The space of $n\times n$ SPD matrices, when provided with some geometric structures like a Riemannian metric, forms SPD manifold $\operatorname{Sym}_{n}^{+}$ (arsigny:inria-00070423). Data lying on SPD manifolds are commonly encountered in various domains (HuangGool17; NguyenHG2018; BrooksRieBatNorm19; NguyenGauLocalPool19; Nguyen_2021_ICCV; SukthankerSPDNetNAS21; NguyenNeurIPS22; NguyenECCV22; NguyenGyroMatMans23). In many applications, the use of Euclidean calculus on SPD manifolds often leads to unsatisfactory results (arsigny:inria-00070423). To tackle this issue, many Riemannian structures for SPD manifolds have been introduced. In this work, we focus on two widely used Riemannian metrics, i.e., Affine-Invariant (AI) (pennec:inria-00070743) and Log-Euclidean (LE) (arsigny:inria-00070423) metrics, and a recently introduced Riemannian metrics, i.e. Log-Cholesky (LC) metrics (Lin_2019) that offer some advantages over Affine-Invariant and Log-Euclidean metrics.
@@ 2.2 Grassmann Manifolds
Grassmann manifolds $\operatorname{Gr}_{n,p}$ are the collection of linear subspaces of fixed dimension $p$ of the Euclidean space $\mathbb{R}^{n}$ (Edelman98). Data lying on Grassmann manifolds arise naturally in many applications (Absil2007; bendokat2020grassmann). Points on Grassmann manifolds can be represented from different perspectives (bendokat2020grassmann). Two typical approaches use projection matrices or those with orthonormal columns. Each of them can be effective in some problems but might be inappropriate in some other contexts (NguyenNeurIPS22). Although geometrical descriptions of Grassmann manifolds have been given in numerous works (Edelman98), some computational issues remain to be addressed. For instance, the question of how to effectively perform backpropagation with the Grassmann logarithmic map in the projector perspective remains open.
@@ 2.3.1 Neural Networks on SPD Manifolds
The work in HuangGool17 introduces SPDNet with three novel layers, i.e., Bimap, LogEig, and ReEig layers that has become one of the most successful architectures in the field. In BrooksRieBatNorm19, the authors further improve SPDNet by developing Riemannian versions of batch normalization layers. Following these works, some works (NguyenHandRecgCVPR19; Nguyen_2021_ICCV; WangSymNet21; kobler2022spdbatchnorm; Ju_2023) design variants of Bimap and batch normalization layers in SPD neural networks. The work in ChakrabortyManifoldNet20 presents a different approach based on intrinsic operations on SPD manifolds. Their proposed layers have nice theoretical properties. A common limitation of the above works is that they do not provide necessary mathematical tools for constructing many essential building blocks of DNNs on SPD manifolds. Recently, some works (NguyenECCV22; NguyenNeurIPS22; NguyenGyroMatMans23) take a gyrovector space approach that enables natural generalizations of some building blocks of DNNs, e.g., MLR for SPD neural networks.
@@ 2.3.2 Neural Networks on Grassmann Manifolds
In HuangAAAI18, the authors propose GrNet that explores the same rule of matrix backpropagation (Ionescu2015) as SPDNet. Some existing works (WangGrasNetNPL20; SouzaCVPRW20) are also inspired by GrNet. Like their SPD counterparts, most existing Grassmann neural networks are not built upon a mathematical framework that allows one to generalize a broad class of DNNs to Grassmann manifolds. Using a gyrovector space approach, NguyenGyroMatMans23 has shown that some concepts in Euclidean spaces can be naturally extended to Grassmann manifolds.
@@ 3.1 Notation
Let $\mathcal{M}$ be a homogeneous Riemannian manifold, $T_{\mathbf{P}}\mathcal{M}$ be the tangent space of $\mathcal{M}$ at $\mathbf{P}\in\mathcal{M}$. Denote by $\exp(\mathbf{P})$ and $\log(\mathbf{P})$ the usual matrix exponential and logarithm of $\mathbf{P}$, $\operatorname{Exp}_{\mathbf{P}}(\mathbf{W})$ the exponential map at $\mathbf{P}$ that associates to a tangent vector $\mathbf{W}\in T_{\mathbf{P}}\mathcal{M}$ a point of $\mathcal{M}$, $\operatorname{Log}_{\mathbf{P}}(\mathbf{Q})$ the logarithmic map of $\mathbf{Q}\in\mathcal{M}$ at $\mathbf{P}$, $\mathcal{T}_{\mathbf{P}\rightarrow\mathbf{Q}}(\mathbf{W})$ the parallel transport of $\mathbf{W}$ from $\mathbf{P}$ to $\mathbf{Q}$ along geodesics connecting $\mathbf{P}$ and $\mathbf{Q}$. For simplicity of exposition, we will concentrate on real matrices. Denote by $\operatorname{M}_{n,m}$ the space of $n\times m$ matrices, $\operatorname{Sym}^{+}_{n}$ the space of $n\times n$ SPD matrices, $\operatorname{Sym}_{n}$ the space of $n\times n$ symmetric matrices, $\operatorname{S}^{+}_{n,p}$ the space of $n\times n$ SPSD matrices of rank $p\leq n$, $\operatorname{Gr}_{n,p}$ the $p$-dimensional subspaces of $\mathbb{R}^{n}$ in the projector perspective. For clarity of presentation, let $\widetilde{\operatorname{Gr}}_{n,p}$ be the $p$-dimensional subspaces of $\mathbb{R}^{n}$ in the ONB (orthonormal basis) perspective (bendokat2020grassmann). For notations related to SPD manifolds, we use the letter $g\in\{ai,le,lc\}$ as a subscript (superscript) to indicate the considered Riemannian metric, unless otherwise stated. Other notations will be introduced in appropriate paragraphs. Our notations are summarized in Appendix A.
@@ 3.2 Neural Networks on SPD Manifolds
In NguyenECCV22; NguyenNeurIPS22, the author has shown that SPD manifolds with Affine-Invariant, Log-Euclidean, and Log-Cholesky metrics form gyrovector spaces referred to as AI, LE, and LC gyrovector spaces, respectively. We adopt the notations in these works and consider the case where $r=1$ (see NguyenNeurIPS22, Definition 3.1). Let $\oplus_{ai},\oplus_{le}$, and $\oplus_{lc}$ be the binary operations in AI, LE, and LC gyrovector spaces, respectively. Let $\ominus_{ai},\ominus_{le}$, and $\ominus_{lc}$ be the inverse operations in AI, LE, and LC gyrovector spaces, respectively. These operations are given in Appendix G.
@@ 3.2.1 FC Layers in SPD Neural Networks
Our method for generalizing FC layers to the SPD manifold setting relies on a reformulation of SPD hypergyroplanes (NguyenGyroMatMans23). We first recap the definition of SPD hypergyroplanes.
Definition 3.1 (SPD Hypergyroplanes (NguyenGyroMatMans23)).
@@
where $\langle.,.\rangle^{g}_{\mathbf{P}}$ denotes the inner product at $\mathbf{P}$ given by the considered Riemannian metric.
@@
where $\mathbf{I}_{n}$ denotes the $n\times n$ identity matrix, and $\langle.,.\rangle^{g}$ is the SPD inner product in $\operatorname{Sym}_{n}^{+,g}$ (NguyenGyroMatMans23) (see Appendix G.7 for the definition of the SPD inner product).
@@
In DNNs, an FC layer linearly transforms the input in such a way that the $k$-th dimension of the output corresponds to the signed distance from the output to the hyperplane that contains the origin and is orthonormal to the $k$-th axis of the output space. This interpretation has proven useful in generalizing FC layers to the hyperbolic setting (shimizu2021hyperbolic).
@@
Let $(\mathbf{e}_{1},\ldots,\mathbf{e}_{m}),\|\mathbf{e}_{i}\|=1,i=1\ldots,m$ be an orthonormal basis of $\mathbb{R}^{m}$. Let $\langle.,.\rangle^{ai}_{\mathbf{P}}$ be the Affine-Invariant metric computed at $\mathbf{P}\in\operatorname{Sym}_{m}^{+,ai}$ as
@@
Denote by $v_{(i,j)}(\mathbf{X})=\langle\ominus_{ai}\mathbf{P}_{(i,j)}\oplus_{ai}\mathbf{X},\mathbf{W}_{(i,j)}\rangle^{ai},\mathbf{P}_{(i,j)},\mathbf{W}_{(i,j)}\in\operatorname{Sym}^{+,ai}_{n},i\leq j,i,j=1,\ldots,m$. Let $\alpha=\frac{1}{m}(\sqrt{1+m\beta}-1)$. Then the output of an FC layer is computed as $\mathbf{Y}=\exp\big([y_{(i,j)}]_{i,j=1}^{m}\big)$, where $y_{(i,j)}]_{i,j=1}^{m}$ is the matrix having $y_{(i,j)}$ as the element at the $i$-th row and $j$-th column, and $y_{(i,j)}$ is given by
@@
As shown in arsigny:inria-00070423, a Log-Euclidean metric on $\operatorname{Sym}_{n}^{+,le}$ can be obtained from any inner product on $\operatorname{Sym}_{n}$. In this work, we consider a metric that is invariant under all similarity transformations, i.e., the metric $\langle\mathbf{W},\mathbf{V}\rangle^{le}_{\mathbf{I}_{n}}=\operatorname{Tr}(\mathbf{W}\mathbf{V})$. We have the following result.
@@ 3.2.2 Convolutional Layers in SPD Neural Networks
where $\mathbf{x}_{(i,j)}^{l}$ is a receptive field of the $l$-th input channel, $\mathbf{w}^{(l,k)}$ is the filter associated with the $l$-th input channel and the $k$-th output channel, and $b^{k}$ is the bias for the $k$-th output channel. Let $\mathbf{X}_{(i,j)}=\operatorname{concat}(\mathbf{x}_{(i,j)}^{1},\ldots,\mathbf{x}_{(i,j)}^{N_{in}})$, $\mathbf{W}_{k}=\operatorname{concat}(\mathbf{w}^{(1,k)},\ldots,\mathbf{w}^{(N_{in},k)})$, where operation $\operatorname{concat}(.)$ concatenates all of its arguments. Then Eq. (1) can be rewritten (shimizu2021hyperbolic) as
@@
In ChakrabortyManifoldNet20, the authors design a convolution operation for SPD neural networks. However, their method is based on the concept of weighted Fréchet Mean, while ours is built upon the concepts of SPD hypergyroplane and SPD pseudo-gyrodistance from an SPD matrix to an SPD hypergyroplane (NguyenGyroMatMans23). Also, our convolution operation can be used for dimensionality reduction, while theirs always produces an output of the same dimension as the inputs.